# weighted-amenability and percolation

### v1

@@ 2.1. Nonunimodularity and relative weights
If $\Gamma$ acts quasi-transitively, then (4) holds for any $x,y$ in the same $\Gamma$-orbit. Moreover, by definition, in this case there are only finitely many $\Gamma$-orbits, and hence the set $\{m(\Gamma_{x})\mid x\in V\}$ is finite. The converse is also true: if $\{m(\Gamma_{x})\mid x\in V\}$ is finite, then $\Gamma$ is unimodular.
@@ 2.2. Weighted-amenability, examples of nonunimodular graphs
Let $G$ be a graph and $\Gamma$ be a closed quasi-transitive subgroup of $\operatorname{Aut}(G)$. For a vertex $o\in V$ let $r\in\mathbb{N}$ be such that every vertex in $G$ is within distance $r$ of some vertex in $\Gamma o$. Form the graph $G^{\prime}$ from the vertices $\Gamma o$ by joining two vertices by an edge if their distance in $G$ is at most $12r+$ 1. Restriction of the elements of $\Gamma$ to $G^{\prime}$ yields a subgroup $\Gamma^{\prime}\subseteq\operatorname{Aut}(G)$. Then $G^{\prime}$ is connected, $\Gamma^{\prime}$ acts transitively on $G^{\prime}$, and we have the following equivalences:
@@
In the context of nonunimodular transitive graphs one may wonder what happens if we replace finite (resp. infinite) clusters with light (resp. heavy) ones. In Definition 1.3 one can replace the condition ``$\left|F\right|<\infty"$ by ``$\mathbf{w}^{o}(F)<\infty"$ (Lemma 4.2), and “no infinite components” by “no heavy components” in the conclusion of Theorem 2.5. Moreover, by a standard trique one can make site the percolation processes $\mathbf{P}_{\alpha}$ stochastically increasing in $\alpha$, making the property in Theorem 2.5 be the same as hyperfiniteness (see Definition 4.1).
@@
The Diestel–-Leader graph $\operatorname{DL}(k,\ell)$, which was first introduced in [DL01], is constructed in the following way. Let $T_{1}:=T_{1,k}$ and $T_{2}:=T_{1,\ell}$ from Example 2.7. Align the levels of the tress in the opposite directions, i.e., if $x\in T_{1}$ is the $i$-th level (with respect to an arbitrarily chosen root) then its children are in the $i+1$-th level; on the other hand for a vertex $y\in T_{2}$ is the $i$-th level then children are in the $i-1$-th level. Denote by $\mathrm{Level}_{T_{i}}(x)$ the level of a vertex $x$ in $T_{i}$. Define a graph $\operatorname{DL}(k,\ell)$ on the vertex set
@@ 2.3. Percolation thresholds and weighted degrees
Let $\Gamma$ be a closed subgroup of $\operatorname{Aut}(G)$ that acts transitively on $G$ and $\mathbf{w}$ be the induced relative weight function as in (1). Suppose $d$ is the degree in $G$ then
@@
Let $\Gamma$ be a closed subgroup of $\operatorname{Aut}(G)$ that acts transitively on $G$ and $\mathbf{w}$ be the induced relative weight function as in (1). Suppose $d$ is the degree in $G$ then
@@
Theorem 2.16 yields that a configuration $\omega$ contains a heavy cluster with positive probability. Considering such an event, notice that the expected degree in $\omega$ under Bernoulli$(p)$ percolation on it is continuous in $p$. Thus, for $p$ large enough, the inequality (9) is violated, implying that there is a cluster $C\subseteq\omega$ with $p_{h}(C,\Gamma)<1$. $\blacksquare$
@@
Since (2.3) counts the expected outgoing mass from $x$ and (2.3) is an upper bounds on the expected incoming mass to $x$, by TMTP we have that
@@ 2.4. Ends of a graph
Let $\Gamma$ be a closed subgroup of $\operatorname{Aut}(G)$ that acts transitively on $G$ and let $\mathbf{w}$ be the induced relative weight function as in (1). Suppose $\mathbf{P}$ is a $\Gamma$-invariant percolation on $G$, then any cluster that has finitely many vertices of maximal weight is light a.s.
@@ 2.5. Measured group theoretic results
As we mentioned above, some of our motivation came from recent progress in understanding the amenability of mcp (measure class preserving) Borel graphs in measured group theory. So, in this section, we briefly present relevant material. To make this paper as close to self-contained as possible, we provide alternative probabilistic proofs to almost all arguments where we apply measure group theoretic results.
@@
We call a CBER $\mathcal{R}$ on a standard probability space $(X,\mu)$ probability measure preserving (pmp) (resp. measure class preserving or mcp) if for any Borel automorphism $\gamma$ on $X$ that maps every point to an $\mathcal{R}$-equivalent point $\gamma$ preserves $\mu$ (resp. $\mu$-null sets). It follows from the argument of Kechris and Woodin, see [Mil04, Proposition 2.1] or [TZ24, 2.2], that any CBER $\mathcal{R}$ on a probability space $(X,\mu)$ is mcp after discarding a null set. The lack of invariance of mcp CBERs is quantified similarly to Theorem 2.1. This is made precise by the Radon–Nikodym cocycle $(x,y)\mapsto\mathbf{w}^{y}(x):\mathcal{R}\to\mathbb{R}^{+}$ of the orbit equivalence relation $\mathcal{R}$ with respect to the underlying probability measure $\mu$, as defined in [KM04, Section 8]. Here the cocycle $\mathbf{w}^{y}(x)$ can be also interpreted as a relative weight function on $X$, hence we use the same notation as we did for graphs whenever the context is clear. The Radon–Nikodym cocycle “corrects” the failure of invariance of the measure $\mu$, enabling the (tilted) mass transport principle (see [KM04, Section 8] and [Gab05, Page 18]): for each $f:\mathcal{R}\to[0,\infty]$ we have that
$\int\sum_{z\in[x]_{\mathcal{R}}}f(x,z)d\mu(x)=\int\sum_{z\in[y]_{\mathcal{R}}}f(z,y)\mathbf{w}^{y}(z)d\mu(y),$ (13)
where $[x]_{\mathcal{R}}$ denotes the equivalence class of $x$ in $\mathcal{R}$. Like for graphs, where $\mathbf{w}\equiv 1$ case of Theorem 2.1 characterizes unimodular graphs, here the Radon–Nikodym cocycle $\mathbf{w}^{y}(x)\equiv 1$ is exactly when $\mathcal{R}$ is pmp. Hence, just as unimodular transitive graphs turned out to be more approachable, much of the measured group theory has been developed under the assumption of pmp. A big part of this is due to the fact that the theory of cost is developed only for pmp CBERs [Gab98]. The cost of a pmp CBER $\mathcal{R}$ is defined as the infimum of the cost (i.e. half of expected degree) of its graphings (Borel graphs $G$ on $X$ whose connectedness relation is equal to $\mathcal{R}$ a.e.). One can interpret cost is an analog for CBERs of the free rank for groups. Unfortunately, most of the theory of pmp actions and CBERs fails in the mcp setting, for instance, nonamenability of the group does not imply $\mu$-nonamenability of the orbit equivalence relations of its free mcp actions. See [CTT22] and the references therein for further discussion.
@@
A Borel graph $\mathcal{G}$ is called pmp/mcp/$\mu$-amenable/$\mu$-hyperfinite if its connectedness equivalence relation is such. Before the development of the theory of cost, the main approach to studying $\mu$-amenability of pmp graphs was through the number of ends. For instance, Adams [Ada90] showed that an acyclic ergodic pmp graph is $\mu$-amenable exactly when it has $\leqslant 2$ ends a.e. Hence, in the absence of the notion of cost in the mcp setting, it is natural to consider the geometry of the mcp graph, this time together with the behavior of the Radon–-Nikodym cocycle. In [TTD25], Tserunyan and Tucker-Drob generalized the result of Adams to the mcp setting as follows:
@@
Let $\mathcal{G}$ be an acyclic mcp graph and $\mathbf{w}$ be the Radon–Nikodym cocycle of the connectedness relation $\mathcal{R}_{\mathcal{G}}$ with respect to the underlying measure. Then the following are equivalent:
@@
    a.e. $\mathcal{G}$-component has $\leqslant 2$ $\mathbf{w}$-nonvanishing ends,
@@
Moreover, if a.e. $\mathcal{G}$-component has $\leqslant 2$ $\mathbf{w}$-nonvanishing ends then all other ends have only light geodesic paths.
Here $\mathbf{w}$-nonvanishing end and light geodesic paths mean the same as in Subsection 2.4, but with respect to the Radon–Nikodym cocycle.
The main result of [CTT22] is an explicit construction of a $\mu$-nonamenable subforest of $\mathcal{G}$, thereby completing the generalization of the Gaboriau–Ghys result ([Ghy95] and [Gab00, IV.24]) to the mcp setting for locally finite graphs.
@@
Let $\mathcal{G}$ be a locally finite mcp Borel graph and let $\mathbf{w}$ denote its Radon–Nikodym cocycle with respect to the underlying probability measure. If a.e. $\mathcal{G}$-component has $\geqslant 3$ $\mathbf{w}$-nonvanishing ends, then there is a Borel subforest $\mathcal{F}\subseteq\mathcal{G}$ such that for a.e. $\mathcal{F}$-connected component, the space of $\mathbf{w}$-nonvanishing ends of that component is nonempty and perfect (no isolated points). In particular, $\mathcal{G}$ is $\mu$-nonamenable. Moreover, $\mathcal{F}$ can be made ergodic if $\mathcal{G}$ is.
The connection between measured group theory and percolation theory is largely facilitated by the general cluster graphing construction from [Gab05, Section 2.2 and 2.3], which we now recall. Let $G=(V,E)$ be a locally finite countable graph, $o\in V$ be an arbitrarily chosen root vertex, $\Gamma$ be a closed subgroup of $\operatorname{Aut}(G)$ that acts transitively on $G$, and $m$ be its left Haar measure. Note that $\Gamma$ is a totally disconnected locally compact group that might not be discrete. Let $\left(\Omega,\mu\right)$ be the standard Borel space, where we let $\Omega:=\{0,1\}^{E}$ and $\mu$ be a $\Gamma$-invariant percolation process on $G$. First consider the free pmp shift action of $\Gamma$ on it $\theta:\Gamma\curvearrowright\left(\Omega,\mu\right)$, i.e., given a configuration $\omega\in\{0,1\}^{E}$, $\gamma\in\Gamma$, and any two vertices $x,y\in V$ we have $\gamma\omega(x,y)=\omega(\gamma^{-1}x,\gamma^{-1}y)$. The goal is now to pass to a countable subrelation of the orbit equivalence relation of $\theta$, which is done by taking a cross section. Our setting allows us to make such a reduction explicit, for more details in a more general setup we refer the reader to [KKR21, Section 3] and the references therein. Since $G$ is a locally finite graph, the stabilizer of the root $\Gamma_{o}$ is compact and thus we may pick a representative configuration from each set $\{\Gamma_{o}\omega\}_{\omega\in\Omega}$ in a Borel fashion, denote such a collection of representatives by $\widehat{\Omega}\subseteq\Omega$. Then the following subequivalence relation $\widehat{\mathcal{R}}$ of the orbit equivalence relation of $\theta$ is a CBER:
^:=$\widehat{\mathcal{R}}:=\{(\omega,\omega^{\prime})\in\widehat{\Omega}\times\widehat{\Omega}\mid\omega^{\prime}\in\Gamma\omega\},$
where we used that $\Gamma_{o}$ is open (and thus $m(\Gamma_{o})>0$) and the fact that $\Gamma$ is second countable (so that $m$ is a $\sigma$-finite measure) to conclude that $\widehat{\mathcal{R}}$ is countable. We now equip $\widehat{\Omega}$ with a measure that is the push-forward of $\mu$ under the operation of quotienting by $\Gamma_{o}$ as above, i.e., for any measurable set $A\subseteq\widehat{\Omega}$ we define $\widehat{\mu}(A)=\mu(\Gamma_{o}A)$. Notice that by mass transport as in [Gab05, page 18] for any $f:V\times V\times\widehat{\Omega}\to[0,\infty]$
$\int_{\widehat{\Omega}}\sum_{x\in\Gamma o\cap\widehat{\Omega}}f(o,x,\omega)\,d\widehat{\mu}=\int_{\widehat{\Omega}}\sum_{x\in\Gamma o\cap\widehat{\Omega}}f(x,o,\omega)\frac{m(\Gamma_{x})}{m(\Gamma_{o})}\,d\widehat{\mu}.$
In particular, the Radon–Nykodim cocycle of $\widehat{\mathcal{R}}$ with respect to the underlying probability measure exactly coincides with the relative weight function $\mathbf{w}^{o}$ induced by $\Gamma$ on $G$ as in (1) (here we used (4) to connect the weights with the Haar measures of the stabilizers). Finally, we define the cluster graphing $\mathcal{G}^{\mathrm{cl}}$ of $\widehat{\mathcal{R}}$ by placing an edge between $\omega$ and $\gamma\omega$ if and only if $\omega(o,\gamma o)=1$.
The following observation summarizes key properties of the cluster graphing.
@@
Unimodularity and amenability of $\Gamma$ are connected to the respective properties of the cluster connectedness relation $\mathcal{R}^{\mathrm{cl}}(:=\widehat{\mathcal{R}})$ of $\mathcal{G}^{\mathrm{cl}}$ via the following results
@@
    The $\mathcal{G}^{\mathrm{cl}}$-component of a configuration $\omega$ is isomorphic to the $\omega$-cluster of the root.
@@
    The Radon–Nikodym cocycle of $\mathcal{R}^{\mathrm{cl}}$ with respect to $\widehat{\mu}$ is equal to $\mathbf{w}_{\Gamma}$ on $G$, as in (1), evaluated on the corresponding roots. In particular, $\mathcal{R}^{\mathrm{cl}}$ is pmp if and only if $\Gamma$ is unimodular [Gab05, Theorem 2.5].
@@
    By [KKR21, Proposition 6] $\mathcal{R}^{\mathrm{cl}}$ is $\mu$-amenable if and only if $\Gamma$ is amenable.

### v2

@@ 2.1. Nonunimodularity and relative weights
If $\Gamma$ acts quasi-transitively, by definition, there are only finitely many $\Gamma$-orbits, and hence if $\Gamma$ is also unimodular the set $\{\mathbf{w}^{o}(x)\mid x\in V\}$ is finite; the converse is also true: if $\{\mathbf{w}^{o}(x)\mid x\in V\}$ is finite, then $\Gamma$ is unimodular.
@@ 2.2. Weighted-amenability, examples of nonunimodular graphs
Let $G$ be a graph and $\Gamma$ be a closed quasi-transitive subgroup of $\operatorname{Aut}(G)$. For a vertex $o\in V$ let $r\in\mathbb{N}$ be such that every vertex in $G$ is within distance $r$ of some vertex in $\Gamma o$. Form the graph $G^{\prime}$ from the vertices $\Gamma o$ by joining two vertices by an edge if their distance in $G$ is at most $2r+1$. Restriction of the elements of $\Gamma$ to $G^{\prime}$ yields a subgroup $\Gamma^{\prime}\subseteq\operatorname{Aut}(G)$. Then $G^{\prime}$ is connected, $\Gamma^{\prime}$ acts transitively on $G^{\prime}$, and we have the following equivalences:
@@
In the context of nonunimodular transitive graphs one may wonder what happens if we replace finite (resp. infinite) clusters with light (resp. heavy) ones. In Definition 1.3 one can replace the condition $\left|F\right|<\infty$" by $\mathbf{w}^{o}(F)<\infty$" (see Lemma 4.2), and “no infinite components” by “no heavy components” in the conclusion of Theorem 2.5. Moreover, by a standard trick one can make the percolation processes $\mathbf{P}_{\alpha}$ stochastically increasing in $\alpha$, making the property in Theorem 2.5 be the same as hyperfiniteness (see Definition 4.1).
@@
The Diestel–-Leader graph $\operatorname{DL}(k,\ell)$, which was first introduced in [DL01], is constructed in the following way. Let $T_{1}:=T_{1,k}$ and $T_{2}:=T_{1,\ell}$ from Example 2.7. Align the levels of the tress in the opposite directions, i.e., if $x\in T_{1}$ is the $i$-th level (with respect to an arbitrarily chosen root) then its children are in the $i+1$-th level; on the other hand for a vertex $y\in T_{2}$ is the $i$-th level then children are in the i1i-1-th level. Denote by $\mathrm{Level}_{T_{i}}(x)$ the level of a vertex $x$ in $T_{i}$. Define a graph $\operatorname{DL}(k,\ell)$ on the vertex set
@@ 2.3. Percolation thresholds and weighted degrees
Let $\Gamma$ be a closed subgroup of $\operatorname{Aut}(G)$ that acts transitively on $G$ and $\mathbf{w}$ be the induced relative weight function as in (1). Suppose $d$ is the degree in $G$. Then
@@
Let $\Gamma$ be a closed subgroup of $\operatorname{Aut}(G)$ that acts transitively on $G$ and $\mathbf{w}$ be the induced relative weight function as in (1). Suppose $d$ is the degree in $G$. Then
@@
Theorem 2.16 yields that a configuration $\omega$ contains a heavy cluster with positive probability. Considering such an event, notice that the expected degree in $\omega$ under Bernoulli(p)(p) percolation on it is continuous in $p$. Thus, for $p$ large enough, the inequality (9) is violated, implying that there is a cluster $C\subseteq\omega$ with $p_{h}(C,\Gamma)<1$. $\blacksquare$
@@
Since (2.3) counts the expected outgoing mass from $x$ and (2.3) is an upper bounds on the expected incoming mass to $x$, by the TMTP we have that
@@ 2.4. Ends of a graph
Let $\Gamma$ be a closed subgroup of $\operatorname{Aut}(G)$ that acts transitively on $G$ and let $\mathbf{w}$ be the induced relative weight function as in (1). Suppose $\mathbf{P}$ is a $\Gamma$-invariant percolation on $G$. Then any cluster that has finitely many vertices of maximal weight is light a.s.
@@ 2.5. Measured group theoretic results
As we mentioned above, some of our motivation came from recent progress in understanding the amenability of mcp (measure class preserving) Borel graphs in measured group theory. So, in this section, we briefly present relevant material.
@@
We call a CBER $\mathcal{R}$ on a standard probability space $(X,\mu)$ probability measure preserving (pmp) (resp. measure class preserving or mcp) if for any Borel automorphism $\gamma$ on $X$ that maps every point to an $\mathcal{R}$-equivalent point $\gamma$ preserves $\mu$ (resp. $\mu$-null sets). It follows from the argument of Kechris and Woodin, see [Mil04, Proposition 2.1] or [TZ24, 2.2], that any CBER $\mathcal{R}$ on a probability space $(X,\mu)$ is mcp on a $\mu$-conull set. The lack of invariance of mcp CBERs is quantified similarly to Theorem 2.1. This is made precise by the Radon–Nikodym cocycle $(x,y)\mapsto\mathbf{w}_{\nu}^{y}(x):\mathcal{R}\to\mathbb{R}^{+}$ of the orbit equivalence relation $\mathcal{R}$ with respect to the underlying probability measure $\mu$, as defined in [KM04, Section 8]. Here the cocycle $\mathbf{w}_{\nu}^{y}(x)$ can be also interpreted as a relative weight function on $X$, hence we use the same notation as we did for graphs whenever the context is clear. The Radon–Nikodym cocycle “corrects” the failure of invariance of the measure $\mu$, enabling the (tilted) mass transport principle (see [KM04, Section 8] and [Gab05, Page 18]): for each $f:\mathcal{R}\to[0,\infty]$ we have that
$\int\sum_{z\in[x]_{\mathcal{R}}}f(x,z)d\mu(x)=\int\sum_{z\in[y]_{\mathcal{R}}}f(z,y)\mathbf{w}_{\nu}^{y}(z)d\mu(y),$ (13)
where $x]_{\mathcal{R}}$ denotes the equivalence class of $x$ in $\mathcal{R}$. Like for graphs, where $\mathbf{w}_{\nu}\equiv 1$ case of Theorem 2.1 characterizes unimodular graphs, here the Radon–Nikodym cocycle $\mathbf{w}_{\nu}^{y}(x)\equiv 1$ is exactly when $\mathcal{R}$ is pmp. Hence, just as unimodular transitive graphs turned out to be more approachable, much of the measured group theory has been developed under the assumption of pmp. A big part of this is due to the fact that the theory of cost is developed only for pmp CBERs [Gab98]. The cost of a pmp CBER $\mathcal{R}$ is defined as the infimum of the cost (i.e. half of expected degree) of its graphings (Borel graphs $G$ on $X$ whose connectedness relation is equal to $\mathcal{R}$ a.e.). One can interpret cost is an analog for CBERs of the free rank for groups. Unfortunately, most of the theory of pmp actions and CBERs fails in the mcp setting, for instance, nonamenability of the group does not imply $\mu$-nonamenability of the orbit equivalence relations of its free mcp actions. See [CTT22] and the references therein for further discussion.
@@
[…]
@@
Let $\mathcal{G}$ be an acyclic mcp graph on a standard probability space $(X,\mu)$ and $\mathbf{w}_{\mu}$ be the Radon–Nikodym cocycle of the connectedness relation $\mathcal{R}_{\mathcal{G}}$ with respect to $\mu$. Then the following are equivalent:
@@
    a.e. $\mathcal{G}$-component has $\leqslant 2$ $\mathbf{w}_{\mu}$-nonvanishing ends,
@@
Moreover, if a.e. $\mathcal{G}$-component has $\leqslant 2$ $\mathbf{w}_{\mu}$-nonvanishing ends then all other ends have only light geodesic paths.
Here $\mathbf{w}_{\mu}$-nonvanishing end and light geodesic paths mean the same as in Subsection 2.4, but with respect to the Radon–Nikodym cocycle $\mathbf{w}_{\mu}$.
The main result of [CTT22] is an explicit construction of a $\mu$-nonamenable subforest of $\mathcal{G}$, which generalizes a theorem of Gaboriau–Ghys ([Ghy95] and [Gab00, IV.24]) to the mcp setting for locally finite graphs.
@@
Let $\mathcal{G}$ be a locally finite mcp graph on a standard probability space $(X,\mu)$ and $\mathbf{w}_{\mu}$ be the Radon–Nikodym cocycle of the connectedness relation $\mathcal{R}_{\mathcal{G}}$ with respect to $\mu$. If a.e. $\mathcal{G}$-component has $\geqslant 3$ $\mathbf{w}_{\mu}$-nonvanishing ends, then there is a Borel subforest $\mathcal{F}\subseteq\mathcal{G}$ such that for a.e. $\mathcal{F}$-connected component, the space of $\mathbf{w}_{\mu}$-nonvanishing ends of that component is nonempty and perfect (no isolated points). In particular, $\mathcal{G}$ is $\mu$-nonamenable. Moreover, $\mathcal{F}$ can be made ergodic if $\mathcal{G}$ is.
The connection between measured group theory and percolation theory is largely facilitated by the general cluster graphing construction [Gab05, Section 2.2 and 2.3], see [CTT22, Section 5B] for a presentation tailored to the nonunimodular/mcp settings. Here we briefly summarize the general construction. Observation 2.23 summarizes all the properties of the construction we will use, hence the reader may continue from there.
Let $G=(V,E)$ be a locally finite connected rooted graph on a vertex set $V$ with root $o\in V$, $\Gamma\subseteq\operatorname{Aut}(G)$ be a closed subgroup, and $\mathbf{P}$ be a $\Gamma$-invariant bond percolation on $G$. In the case when $\Gamma$ is a countable finitely generated group and $G$ is its Cayley graph with respect to a symmetric set of generators $\Sigma$, the cluster graphing can be defined in a straightforward way ([GL09, Pet17]) as follows. Consider the shift action of $\Gamma$ on $2^{E}$ and the corresponding Schreier graph $\mathcal{G}$, namely place an edge between $\omega,\omega^{\prime}\in 2^{E}$ whenever there exists $\gamma\in\Sigma$ such that $\gamma\omega=\omega^{\prime}$. Then the cluster graphing $\mathcal{G}^{\mathrm{cl}}$ is defined as a subgraphing of $\mathcal{G}$ where the edge $(\omega,\omega^{\prime})\in\mathcal{G}$ is retained if and only if $(o,\gamma o)\in\omega$. Notice that this construction is $\Gamma$-invariant, i.e., in this case $\gamma^{-1}\omega^{\prime}=\omega$ and $(o,\gamma o)\in\omega$ if and only if $(o,\gamma^{-1}o)\in\omega^{\prime}$.
For a general closed group $\Gamma\subseteq\operatorname{Aut}(G)$ the definition of a cluster graphing requires a few technical steps. Formally one considers an ergodic free pmp action $\Gamma$ on a standard probability space $(X,\mu)$ that factors onto $(2^{E},\mathbf{P})$ (i.e., admits a $\Gamma$-equivariant projection $\pi:X\to 2^{E}$ such that $\pi_{*}(\mu)=\mathbf{P}$) and its extension to the diagonal action of $\Gamma$ on $X\times V$ (we think of it as assigning root vertices to elements of $X$). The cluster graphing is then defined on a standard probability space $(Y,\nu)$, where $Y=(X\times V)/\Gamma$. Fixing an $o\in V$, for each element of $Y$ we can take the representative of the form $(x,o)$, and define probability measure $\nu$ on $Y$ as $\nu(A):=\mu(\{x\in X,[(x,o)]_{\Gamma}\in A\})$. Now, the action of $\Gamma$ on $Y$ may not preserve $\nu$ (as is the case when $\Gamma$ in nonunimodular), but it is always mcp on a $\nu$-conull set (as mentioned above e.g. by [Mil04, Proposition 2.1]). Similarly to the above, we first define a graphing $\mathcal{G}$ whose connected components are isomorphic to $G$ (also known as the graphing of the rerooting relation): by setting two elements $\widetilde{x}:=[(x,u)]_{\Gamma}$ and $\widetilde{y}:=[(y,v)]_{\Gamma}$ from $Y$ to be adjacent in $\mathcal{G}$ if and only if there exists $\gamma\in\Gamma$ so that $x=\gamma y$ and $(u,\gamma v)\in E$. That is, one can choose different representatives so that their first coordinates are the same, and their second coordinates are connected by an edge in $G$. Note that since $\pi$ is $\Gamma$-equivariant it is well defined on $Y$ (for any $\widetilde{x}=[(x,u)]_{\Gamma}$ we have $\pi(\widetilde{x})=\pi(x)$). We now define the cluster graphing $\mathcal{G}^{\mathrm{cl}}\subseteq\mathcal{G}$ by retaining the edges $([(x,u)]_{\Gamma},[(x,v)]_{\Gamma})$ of $\mathcal{G}$ such that $(u,v)\in\pi(x)$. Thus, the $\mathcal{G}^{\mathrm{cl}}$-component of a point $\widetilde{x}=[(x,o)]_{\Gamma}\in Y$ is isomorphic to the cluster of the root $o$ in the percolation configuration $\pi(\widetilde{x})=\pi(x)$. Moreover, the Radon–Nikodym cocycle of $\mathcal{G}^{\mathrm{cl}}$ with respect to $\nu$ is given by the relative weight function $\mathbf{w}_{\Gamma}$ on $G$ induced by $\Gamma$ (as defined in (1)), see [CTT22, Lemma 5.7] for the proof.
The following observation summarizes its key properties.
@@
Amenability of a closed group $\Gamma\subseteq\operatorname{Aut}(G)$ and the properties of a $\Gamma$-invariant percolation $\mathbf{P}$ on $G$ are connected to the properties of the cluster graphing $\mathcal{G}^{\mathrm{cl}}$ via the following results.
@@
    The $\mathcal{G}^{\mathrm{cl}}$-component of $\widetilde{x}=[(x,o)]_{\Gamma}\in Y$ is isomorphic to the cluster of $o$ in $\pi(x)$.
@@
    The Radon–Nikodym cocycle $\mathbf{w}_{\nu}$ of $\mathcal{G}^{\mathrm{cl}}$ with respect to $\nu$ is induced by $\mathbf{w}_{\Gamma}$ (as in (1))
    $\mathbf{w}_{\nu}([x,v]_{\Gamma},[x,u]_{\Gamma}):=\mathbf{w}_{\nu}^{[x,u]_{\Gamma}}([x,v]_{\Gamma}):=\mathbf{w}_{\Gamma}^{u}(v),$ (14)
@@
    If $\mathcal{G}^{\mathrm{cl}}$ is nowhere amenable then $\Gamma$ is nonamenable. This follows from [KKR21, Proposition 6] and [JKL02, Page 49].